At today's spot exchange rates 1 U.S. dollar can be exchanged for 10 Mexican pesos or for 110.38 Japanese yen. You have pesos that you would like to exchange for yen. What is the cross rate between the yen and the peso; that is, how many yen would you receive for every peso exchanged? Round your answer to two decimal places.
Answer #1

Cross rate between the yen and the peso = (Yen/U.S. Dollar exchange rate)/(Peso/U.S. Dollar exchange rate)

Cross rate between the yen and the peso = 110.38/10 = 11.04 yen/peso

i.e. For every one peso, you will receive 11.04 Yen.